Brianna borrows $500 on January 1,2015. She repays the loand with 20 annual payments starting January 1, 2016. The payments in even numbered years are y each and the oayments in odd numbered years are x each. If i=.08 and the total of all 20 loan payments is 1,023, find x and y.
Answer #1
Sum of discount factors
of odd years
5.10
Sum of discount factors
of even years
4.72

10x+10y= 1023

Hence x= (1023-10y)/10

x= 102.3-y

PV of annual payments = 500 = 5.1x+4.72y

500= 5.1*(102.3-y)+4.72y

500 = 521.73-5.1y+4.72y

0.38y=21.73

y= $57.18

x= 102.3-57.18 = 45.12

Workings

Odd Years Discount factor Even Years Discount factor
1 0.93 2 0.86
3 0.79 4 0.74
5 0.68 6 0.63
7 0.58 8 0.54
9 0.50 10 0.46
11 0.43 12 0.40
13 0.37 14 0.34
15 0.32 16 0.29
17 0.27 18 0.25
19 0.23 20 0.21
Sum of discount factors
of odd years
5.10 Sum of discount factors
of even years
4.72

Book2 - Excel Sign in - . AutoSave off 2 File Home H D . Draw Page Layout Insert Formulas Data Review View Help Tell me what